# Multimodal Conversation Structure Understanding

### v2

@@ Abstract
Conversations are usually structured by roles—who is speaking, who’s being addressed, and who’s listening—and unfold in threads that break with changes in speaker floor or topical focus. While large language models (LLMs) have shown incredible capabilities in dialogue and reasoning, their ability to understand fine-grained conversational structure, especially in multi-modal, multi-party settings, remains underexplored. To address this gap, we introduce a suite of tasks focused on conversational role attribution (speaker, addressees, side-participants) and conversation threading (utterance linking and clustering), drawing on conversation analysis and sociolinguistics. To support those tasks, we present a human annotated dataset of 4,398 annotations for speakers and reply-to relationship, 5,755 addressees, and 3,142 side-participants.
We evaluate popular audio-visual LLMs and vision–language models on our dataset, and our experimental results suggest that multimodal conversational structure understanding remains challenging. The most performant audio-visual LLM outperforms all vision–language models across all metrics, especially in speaker and addressee recognition. However, its performance drops significantly when conversation participants are anonymized. The number of conversation participants in a clip is the strongest negative predictor of role-attribution performance, while acoustic clarity (measured by pitch and spectral centroid) and detected face coverage yield positive associations. We hope this work lays the groundwork for future evaluation and development of multimodal LLMs that can reason more effectively about conversation structure.
@@ 1 Introduction
Multi-party conversation understanding involves identifying and structuring interactions between multiple speakers and recipients in a face-to-face setting. Such conversations typically involve complex floor-claiming and turn-taking patterns marked by various pragmatic cues. Modeling these structures is useful for any dialogue system, and it has the potential to shed light on the power dynamics between interlocutors, reveal implicit social hierarchies, and enhance our understanding of how conversational roles shift over time, among other things. While multimodal large language models (LLMs) have shown potential in reasoning over complex videos and dialogues, whether they can adequately parse the structure of conversation—resolve the reply-to relationship between utterances, or attribute roles like speakers and addressees—remains underexplored. To bridge this gap, in this work we make the following contributions:
@@
Figure 1: Our proposed structured prediction task for multimodal conversation structure understanding. Grounded in sociolinguistic and conversational analysis [1, 2, 3, 4, 5], the task requires predicting, for each utterance in the given clip: the speaker, addressee(s), side-participants, and the utterance it replies to. The example, taken from the Big Bang Theory in TVQA [28, 29], illustrates our unified formulation, which treats conversational role attribution and conversation disentanglement as complementary subtasks for modeling the interactional dynamics of dialogue. Further analysis of this example can be found in §A.1.3.
  • We draw on literature in sociolinguistics [1, 2, 3] and conversational organization [4, 5] to devise a framework for multimodal multi-party conversation structure understanding that seeks to capture the interactional patterns of interlocuters. Building on this framework, we formally introduce a suite of tasks (Fig. 1) focused on conversational role attribution (speaker, addressees, side-participants) and conversation disentanglement (utterance linking and clustering).

### v3

@@ Abstract
While multimodal large language models (LLMs) excel at dialogue, whether they can adequately parse the structure of conversation—conversational roles and threading—remains underexplored. In this work, we introduce a suite of tasks and release TV-MMPC, a new annotated dataset, for multimodal conversation structure understanding. Our evaluation reveals that while all multimodal LLMs outperform our heuristic baseline, even the best-performing model we consider experiences a substantial drop in performance when character identities of the conversation are anonymized. Beyond evaluation, we carry out a sociolinguistic analysis of 350,842 utterances in TVQA. We find that while female characters initiate conversations at rates in proportion to their speaking time, they are 1.2 times more likely than men to be cast as an addressee or side-participant, and the presence of side-participants shifts the conversational register from personal to social.
Multimodal Conversation Structure Understanding
Kent K. Chang    Mackenzie Hanh Cramer    Anna Ho Ti Ti Nguyen   Yilin Yuan   David Bamman University of California, Berkeley kentkchang@berkeley.edu
@@ 1 Introduction
Figure 1: Our proposed structured prediction task for multimodal conversation structure understanding. Grounded in sociolinguistic and conversation analysis (Goffman, 1981, 1983; Ng and Bradac, 1993; Goodwin, 1981; Clark and Carlson, 1982), the task requires predicting, for each utterance in the given clip: the speaker, addressee(s), side-participants, and the utterance it replies to. The example, taken from The Big Bang Theory in TVQA Lei et al. (2018, 2020), illustrates our unified formulation, which treats conversational role attribution and conversation disentanglement as complementary subtasks for modeling the interactional dynamics of dialogue. Further analysis of this example can be found in Appendix §F.3.
Multi-party conversation understanding involves identifying and structuring interactions between multiple speakers and recipients in a face-to-face setting Akhtiamov et al. (2019); Lerner et al. (2022); Korbar et al. (2024). Such conversations typically involve complex floor-claiming Allan (2002); Elsner and Charniak (2008, 2010); Wang et al. (2020) and turn-taking patterns Sacks et al. (1974); Hawes et al. (2009); Zhou et al. (2018) marked by various pragmatic cues, which is a challenging problem in NLP Gu et al. (2022); Tan et al. (2023); Penzo et al. (2024). Modeling these interactions—often described in sociolinguistics as ritualized social encounters Goffman (1981)—further enables us to shed light on power dynamics and reveal implicit social hierarchies in conversation.
While multimodal large language models (LLMs) have shown potential in reasoning over complex videos and dialogues Song et al. (2023); Teng et al. (2023); Zhang et al. (2023); Bhattacharyya et al. (2023); Sun et al. (2024), whether they can adequately parse the structure of conversation—resolve the reply-to relationship between utterances, or attribute roles like speakers and addressees—remains underexplored Le Minh et al. (2018); Inoue et al. (2025). To bridge this gap:
  • We draw on sociolinguistics (Goffman, 1981, 1983; Goodwin, 1981; Clark and Carlson, 1982; Ng and Bradac, 1993) to devise a framework and introduce a suite of tasks (Fig. 1) for capturing the interactional patterns of interlocutors in the multimodal context.